OpenStudy (diyadiya):

If the equation \(x^2+ax+b=0\) & \(x^2+bx+a=0\) have common root then the numerical value of a+b is

OpenStudy (slaaibak):

x^2 + ax + b = x^2 + bx + a ax - bx + b - a = 0 a(x - 1) - b(x - 1) = 0 (a - b)(x - 1) = 0 a - b = 0 therefore a = b Root: x=1 Now sub'ing back: 1 + a + b = 0 a + b = -1

OpenStudy (diyadiya):

Wow Thankyou =D

OpenStudy (slaaibak):

Not sure if this is correct. But I hope it helps

OpenStudy (diyadiya):

Why x=1?

OpenStudy (slaaibak):

(a - b)(x - 1) = 0 Therefore x must be equal to 1 to satisfy the equation.

OpenStudy (diyadiya):

Btw answer is Right :) i have the final answer =D

OpenStudy (diyadiya):

Ohhhh yeahhhh !!!

OpenStudy (slaaibak):

haha, cool :) Glad I could help!

OpenStudy (diyadiya):

=)